A General Power bond carries a coupon rate of 8.7%, has 9 years until maturity, and sells at a yield to maturity of 7.7%. (Assume annual interest payments.)

a. What interest payments do bondholders receive each year?

b. At what price does the bond sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

c. What will happen to the bond price if the yield to maturity falls to 6.7%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

d. If the yield to maturity falls to 6.7%, will the current yield be less, or more, than the yield to maturity?
Answer #1

a) Interest payments refers to the coupon payments=(Coupon rate)*(Face value)
Suppose the face value of the bond is $1000
Here, the coupon rate is 8.7%
Coupon interest payments=1000*8.7%=$87

b)
Face value=$1000
Coupon payment=$87
Time period=9 years
Yield to maturity=7.7%
We can determine the present value using excel.

The bond sell at a price of $1,063.26

1 Face value $1,000 2 Coupon payment 87 3 Time period 4 Yield to maturity 7.70% 5 Present value= ($1,063.26) 6 Formula used:

c)
If yield to maturity falls to 6.7%
The bond price will change to $1,131.98

1 Face value $1,000 2 Coupon payment 3 Time period 4 Yield to maturity 6.70% 5 Present value= ($1,131.98) 6 Formula used: PV(

d)
Current yield =(Annual coupon payment)/(Current market price of the bond)
When yield to maturity is 7.7%, the current price=$1,063.26, the current yield =87/1063.26=0.081823825 or 8.18% (Rounded to two decimal places)
When yield to maturity is 6.7%, the current price=$1,131.98, the current yield =87/1131.98=0.076856482 or 7.69% (Rounded to two decimal places)

So, if yield to maturity falls to 6.7%, the current yield will be less.